OpenStudy (anonymous):

@ganeshie8 @jhonyy9 @satellite73 I need help, please help me.

OpenStudy (anonymous):

A system of linear inequalities is shown below. x - y > 3 y + x ≤ 2 Describe the steps to graph the solution set to the system of inequalities.

OpenStudy (anonymous):

x - y > 3 y + x <_ 2

OpenStudy (imstuck):

Do one of the inequalities at a time.

OpenStudy (imstuck):

I can help you if you'd like. I teach this to my algebra students in a way that seems to make it quite a bit easier for them in the long run.

OpenStudy (anonymous):

Then find the intersection of 2 graphs combined in one graph. That is the answer.

OpenStudy (anonymous):

yes plz teach me @IMStuck

OpenStudy (anonymous):

hello @IMStuck

OpenStudy (anonymous):

@WilliamZhang can you help me

OpenStudy (anonymous):

what do u want to know

OpenStudy (anonymous):

how to work this out

OpenStudy (anonymous):

read my reply before, I have done for you.

OpenStudy (anonymous):

U dont understand what you did on the bottom

OpenStudy (anonymous):

I *

OpenStudy (anonymous):

The shadowed space is the area presenting the inequality for y. Draw 2 graphs in one paper and find the intersection of the shadowed space. That is your answer .

OpenStudy (anonymous):

can you show me how

OpenStudy (anonymous):

For the question i say x-y>3 y<x-3 everything below y=x-3 is shaded

OpenStudy (anonymous):

y+x<_2 y<_-x+2 everything is shaded below y=-x+2

OpenStudy (anonymous):

x - y > 3 y + x ≤ 2 is your inequality system. Any possible y has to fit in this requirement, which means by the graphs, any y in the shaded space in the intersection space fits the system.

OpenStudy (anonymous):

what do you mean by any y in the shaded space, im really confused on how this is set up.

OpenStudy (anonymous):

How do i set this up to fit exactly what this question is asking?

OpenStudy (anonymous):

i get that part but it doesnt match what i have

OpenStudy (anonymous):

For the first equation the point would be (3,0) it would be a dotted line shading to the right.

OpenStudy (anonymous):

what do you have, graph it or write it down.

OpenStudy (anonymous):